# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v1

@@ 6.2. Design Recommendations to Support Authentic Self-Presentation Online
Informed by data derived from our study and SMSM, we propose design recommendations for fostering authentic self-presentation on social media.
@@
Provide air cover for audience management. Creating a safe space is essential for facilitating automatic self-presentation. While Automatic processes may seem less relevant to authenticity on social media, it is important to recognize that Controlled self-presentations require significant effort and can be mentally exhausting, potentially leading to self-regulation challenges. This exhaustion not only increases problematic social media use but may also diminish users’ effectiveness in managing their subsequent self-presentations, such as oversharing or appearing arrogant. By protecting the cognitive resources of users, particularly children, we can enhance their capacity for more accurate and authentic self-presentation.
@@ 6.3. Limitations and Future Work
Our study has several limitations. The first is that our participant sample includes significantly more people identifying as girls than as boys or non-binary. Further, BeReal introduced many new features since our data collection: (1) the integration of Spotify, where the cover art of the song or podcast a BeReal user is listening to will appear in the bottom right corner of their post, (2) a bonus BeReal where BeReal allows users to post an additional two BeReals throughout the day — but only if they post their first during the allotted two-minute window, (3) branded accounts (e.g., Adidas, Notion), and (4) “Groups” which resemble group chats, to name a few. A lot of how the platform norm is perceived is likely to have changed with the introduction of the new features. It would be useful for future work to explore how this extra disclosure and extra element of gamification support or undermine healthy and desired forms of authentic self-presentation or create further privacy concerns or pressure.
Future studies might also consider exploring methods to support more comprehensive authentic sharing integral to adolescent relationship-building. Specifically, while BeReal’s design indeed offers a new approach to social media—encouraging users to share more authentically—our study’s participants highlighted that the platform focuses primarily on one facet of self-disclosure: transparent photo sharing. They also mentioned the presence of potentially harmful, undesired, or incomplete forms of authenticity on BeReal. As a result, there remains a need for future research to explore novel designs that could cultivate different modes of sharing that transcend a single form of self-disclosure or self-presentation on social media, moving towards, for example, more intimate self-disclosure or deliberate self-presentation that our participants expressed a desire for.
@@ 7. Conclusion
In this paper, we explore the impact of design on promoting authenticity in social media. We conducted a qualitative interview study with 29 participants who use BeReal, a social media platform created to reduce self-presentation pressures and inauthenticity that arise from excessive content curation and strategic self-presentation. Our findings demonstrate that the affordances of BeReal effectively generate a space for authentic self-presentation, where users share their unfiltered lives and engage in relatively little social comparison. However, we also find that transparent photo sharing is only a single facet of our teen participants’ understanding of authentic self-presentation, and young people are eager for online spaces that allow them to share more than just the details of their day. This study contributes to the important discourse about the influence of social media on teen relationship-building and presents a framework and guideline for developing platforms that support the different relationship and self-presentation needs of the youth.
@@ Acknowledgements.
- Carpenter and Greene (2015a) Amanda Carpenter and Kathryn Greene. 2015a. Social Penetration Theory. In The International Encyclopedia of Interpersonal Communication (1 ed.), Charles R Berger, Michael E Roloff, Steve R Wilson, James Price Dillard, John Caughlin, and Denise Solomon (Eds.). Wiley, 1–4. https://doi.org/10.1002/9781118540190.wbeic160
- Carpenter and Greene (2015b) Amanda Carpenter and Kathryn Greene. 2015b. Social penetration theory. The international encyclopedia of interpersonal communication (2015), 1–4.
@@
- Naegele and Goffman (1956) Kaspar D Naegele and Erving Goffman. 1956. The presentation of self in everyday life. Am. Sociol. Rev. 21, 5 (Oct. 1956), 631.

### v2

@@ 6.2. Design Recommendations to Support Authentic Self-Presentation Online
Informed by data derived from our study and the SMSM, we propose design recommendations for fostering authentic self-presentation on social media.
@@
Provide air cover for audience management. Creating a safe space is essential for facilitating Automatic self-presentation. While automatic processes may seem less relevant to authenticity on social media, it is important to recognize that Controlled self-presentations require significant effort and can be mentally exhausting, potentially leading to self-regulation challenges. This exhaustion not only increases problematic social media use but may also diminish users’ effectiveness in managing their subsequent self-presentation, such as oversharing. By protecting the cognitive resources of users, particularly children, we can enhance their capacity for more accurate and authentic self-presentation.
@@ 6.3. Limitations and Future Work
Our study has several limitations. The first is that our participant sample includes significantly more people identifying as girls than as boys or non-binary. Further, BeReal introduced many new features since our data collection: (1) the integration of Spotify, where the cover art of the song or podcast a BeReal user is listening to will appear in the bottom right corner of their post, (2) a bonus BeReal where BeReal allows users to post an additional two BeReals throughout the day—but only if they post their first during the allotted two-minute window, (3) branded accounts (e.g., Adidas, Notion), and (4) “Groups” which resemble group chats, to name a few. Perceptions of platform norms have likely to have changed with the introduction of the new features. It would be useful for future work to explore how this extra disclosure and extra element of gamification support or undermine healthy and desired forms of authentic self-presentation or create further privacy concerns or pressure.
Future studies might also consider exploring methods to support more comprehensive authentic sharing integral to adolescent relationship-building. Specifically, while BeReal’s design indeed offers a new approach to social media—encouraging users to share more authentically—our study’s participants highlighted that the platform focuses primarily on one facet of self-disclosure: transparent photo sharing. They also mentioned the presence of potentially harmful, undesired, or incomplete forms of “authentic” self-presentation on BeReal. As a result, there remains a need for future research to explore novel designs that could cultivate different modes of sharing that transcend a single form of self-disclosure or self-presentation on social media, moving towards, for example, more intimate self-disclosure or deliberate self-presentation that our participants expressed a desire for.
@@ 7. Conclusion
In this paper, we explore the impact of design on promoting authenticity in social media. We conducted a qualitative interview study with 29 teen participants who use BeReal, a social media platform created to reduce self-presentation pressures and perceived lack of authenticity that arise from excessive content curation and strategic self-presentation. Our findings demonstrate that the affordances of BeReal effectively generate a space for authentic self-presentation, where users share their unfiltered lives and engage in relatively little social comparison. However, we also find that transparent photo sharing is only a single facet of our teen participants’ understanding of authentic self-presentation, and young people are eager for online spaces that allow them to share more than just the details of their day. This study contributes to the important discourse about the influence of social media on teen relationship-building and presents a framework and design guidance for developing platforms that support the different relationship and self-presentation needs of the youth.
@@ Acknowledgements.
- Altman and Taylor (1973) Irwin Altman and Dalmas A Taylor. 1973. Social penetration: The development of interpersonal relationships. Holt, Rinehart & Winston.
@@
- Canady (2023) Valerie A Canady. 2023. APA calls for improved social media messaging to parents, teens. Mental Health Weekly 33, 8 (2023), 7–8.
- Carpenter and Greene (2015) Amanda Carpenter and Kathryn Greene. 2015. Social Penetration Theory. In The International Encyclopedia of Interpersonal Communication (1 ed.), Charles R Berger, Michael E Roloff, Steve R Wilson, James Price Dillard, John Caughlin, and Denise Solomon (Eds.). Wiley, 1–4. https://doi.org/10.1002/9781118540190.wbeic160
@@
- TikTok (2022) TikTok. 2022. Introducing more ways to create and connect with TikTok Now. https://newsroom.tiktok.com/en-us/introducing-tiktok-now. Accessed: 2024-8-25.
@@
- Turkle (2016) Sherry Turkle. 2016. Reclaiming conversation: The power of talk in a digital age. Penguin.
@@
- Vazire and Gosling (2004) Simine Vazire and Samuel D Gosling. 2004. e-Perceptions: personality impressions based on personal websites. Journal of personality and social psychology 87, 1 (2004), 123.